The reactant concentration in a first-order reaction was 8.60×10−2 M after 45.0 s and 1.00×10−2 M after 100 s . What is the rate constant for this reaction? Express your answer with the appropriate units.

Indicate the multiplication of units, as necessary, explicitly either with a multiplication dot or a dash. Please provide units.
Answer #1

we have:

[A]o = 0.086 M

[A] = 0.010 M

t = t2-t1 = 100 - 45.0 = 55.0 s

use integrated rate law for 1st order reaction

ln[A] = ln[A]o - k*t

ln(1*10^-2) = ln(8.6*10^-2) - k*55

-4.605 = -2.453 - k*55

k*55 = 2.152

k = 3.912*10^-2 s-1

Answer: 3.91*10^-2 s-1